OpenStudy (anonymous):

what two positive real numbers whose product is 50 have the smallest possible sum?

jimthompson5910 (jim_thompson5910):

Let x and y be the two numbers Also let S be the sum of the two numbers. So S = x+y their product is 50 which means x*y = 50 ---> y = 50/x

OpenStudy (anonymous):

ok got that much

jimthompson5910 (jim_thompson5910):

so we can say S = x+y S = x + x/50 ... plug in y = 50/x now your task is to minimize S given that S = x + x/50

jimthompson5910 (jim_thompson5910):

oops i meant to say 50/x, not x/50

jimthompson5910 (jim_thompson5910):

so S = x + 50/x

OpenStudy (anonymous):

w/e i got u

jimthompson5910 (jim_thompson5910):

ok good

OpenStudy (anonymous):

where do i go from here though?

jimthompson5910 (jim_thompson5910):

graph x + 50/x and use the min point feature to find the minimum point

OpenStudy (anonymous):

but what if i just wanted to use calculus

jimthompson5910 (jim_thompson5910):

you would derive with respect to x, set the derivative equal to zero, then solve for x

jimthompson5910 (jim_thompson5910):

once you have x, you can find y

OpenStudy (anonymous):

derivative of which one?

jimthompson5910 (jim_thompson5910):

S = x + 50/x

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

derive to get S ' = 1 - 50/x^2

OpenStudy (anonymous):

thanks so much

jimthompson5910 (jim_thompson5910):

np